OpenStudy (anonymous):

Select the choice below that could be the first step in solving the following system of equations using substitution:8x + 2y = 115y − x + 12 = 0 Answer Multiply the first equation by -4 Multiply the second equation by 2 Write the second equation in terms of x Replace the right side of the first equation with the second equation

OpenStudy (callisto):

@greener6 What's your thought on this question?

OpenStudy (anonymous):

i dont now it confuses me on how to do it

OpenStudy (anonymous):

Write the second equation in terms of x first bring in form of 1st equation like x terms first then y terms....

OpenStudy (callisto):

Split 8x + 2y = 115y − x + 12 = 0 into two questions for better observation. 8x + 2y = 0 115y − x + 12 = 0 Which method would you choose to solve these equations - substitution or elimination?

OpenStudy (anonymous):

substitution